OpenStudy (k12awesomeness):

Will Medal and Fan!!!! Check my answer please!! Which equation is represented by the table? x y (x, y) 0 2 (0, 2) 1 –1 (1, –1) 2 –4 (2, –4) A. y = x – 1 B. y = 2x C. y = –3x + 2 D. y = –2x I think C maybe

OpenStudy (k12awesomeness):

@FaiqRaees @scoobydo @Hauntedwoodsgal @Ms-Brains @welshfella

OpenStudy (k12awesomeness):

@Av0Cad0

OpenStudy (hauntedwoodsgal):

What do you think

OpenStudy (k12awesomeness):

C maybe

OpenStudy (misssmartiez):

Plug it in.

OpenStudy (k12awesomeness):

huh?

OpenStudy (misssmartiez):

0,2. Ok lets just say the function is: –3x + 2

OpenStudy (misssmartiez):

-3 x 0 = ? That plus two is

OpenStudy (welshfella):

plug in x = 0 x = 1 and x= 2 into the equation in C and see if they fit.

OpenStudy (misssmartiez):

-3 x 0 = 0. + 2 is 2. so, 0,2. CORRECT, C is so far correct

OpenStudy (k12awesomeness):

yeah

OpenStudy (misssmartiez):

-3 x 1 = -3. + 2 = -1 (1, –1)

OpenStudy (k12awesomeness):

okay

OpenStudy (misssmartiez):

C, is still so far correct. This is how we do it. (_ This, stands for x. So, we plug whatever shows in this blank, from what it shows us. _ is y. After we finish the function, it should EQUAL what is in here.) (x, y). x = variable isolated. y = solution.

OpenStudy (k12awesomeness):

okay....

OpenStudy (misssmartiez):

Last one: (2, –4) -3 x 2. (which is located in the x place) is -6. + 2? -4. So, 2,-4 is the answer. Therefore, you are correct. C!

OpenStudy (k12awesomeness):

yay!! Thanks everyone!!!

OpenStudy (misssmartiez):

No problem.